OpenStudy (anonymous):

If 3x-10=24, then x=?

OpenStudy (anonymous):

Adding both sides by 10 would be good start

OpenStudy (anonymous):

11 1/3

OpenStudy (hanner_b_nanner):

3x-10=24 +10 +10 3x=34/3 x=11.33333333333333